OpenStudy (anonymous):

FAN AND MEDAL!!!!! PLEASE HELP!!!!! The following function defines a recursive sequence: f(0) = -4 f(1) = 12 f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? A) -4, 12, -28, 60, … B) -4, -12, -28, -60, … C) -4, 12, -18, 54, … D) -4, 12, -18, -54, …

OpenStudy (pandemonium):

B

OpenStudy (anonymous):

B can't be the answer. @Pandemonium

OpenStudy (pandemonium):

no wait let me fix this real quick

OpenStudy (anonymous):

okay

OpenStudy (pandemonium):

i got C but id check with someone else to be sure

OpenStudy (anonymous):

Please explain how you got the answer

OpenStudy (anonymous):

So did you just guess?

OpenStudy (pandemonium):

No my teacher is here helping me and even he's confused what the question is asking but he's trying different methods to find the answer

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

It's florida virtual school, so I wouldn't be surprised if it was entirely wrong.

OpenStudy (anonymous):

@pooja195 please help

OpenStudy (pandemonium):

yea, but me him decide to eliminate the worst possible answers and we got rid of A and B

OpenStudy (anonymous):

okay thx

OpenStudy (pandemonium):

so your left with C and D

OpenStudy (pandemonium):

ok, f(0) = -3 , first term = -3 f(1) = 6, 2nd term = 6 so, as of now our sequence is -3,6 to find nect term, we will use f(n) = -2•f(n -1) - f(n - 2) put n= 2 here, what do u get ?

OpenStudy (anonymous):

-2, but that must be wrong

OpenStudy (pandemonium):

you just replace "n" by "2" there, to get f(2)

OpenStudy (pandemonium):

f(2) = -2 f(1) - f(0) and you already know f(1) =6, f(0) = -3 just plug in values!

OpenStudy (pandemonium):

you need to find f(3) too! f(3) = - 2f(2) - f(1) = - 2(-9) -6 = 18-6 = 12

OpenStudy (anonymous):

That's what I did. Idk how I got such a bogus answer. What did you get?

OpenStudy (pandemonium):

Well A isnt matching and B isnt either

OpenStudy (anonymous):

Can you do it for f(4)? I'd probably get it wrong

OpenStudy (pandemonium):

no that'll probably mess it up but u can try and see if u find a good answer

OpenStudy (anonymous):

okay

OpenStudy (pandemonium):

ok

OpenStudy (anonymous):

I got -13, but it's obviously wrong

OpenStudy (pandemonium):

yea, well my best possible answer would have to be D

OpenStudy (anonymous):

really, why?

OpenStudy (pandemonium):

its the only one close to matching what me and my teacher got but to sure try to find other possible solutions

OpenStudy (anonymous):

ok

OpenStudy (pandemonium):

so yea its either gonna be C or D

OpenStudy (pandemonium):

my best guess would be D but i aint 100% sure